Minden Company introduced a new product last year for which it is trying to find an optimal selling price. Marketing studies suggest that the company can increase sales by 5,000 units for each $2 reduction in the selling price. The company’s present selling price is $91 per unit, and variable expenses are $61 per unit. Fixed expenses are $830,400 per year. The present annual sales volume (at the $91 selling price) is 25,200 units.

Required:

1. What is the present yearly net operating income or loss?

2. What is the present break-even point in unit sales and in dollar sales?

3. Assuming that the marketing studies are correct, what is the maximum annual profit that the company can earn? At how many units and at what selling price per unit would the company generate this profit?

4. What would be the break-even point in unit sales and in dollar sales using the selling price you determined in (3) above (e.g., the selling price at the level of maximum profits)?
Answer #1

1) Present net income (loss)

Sales (25200*91) 2293200
Variable cost (25200*61) 1537200
Contribution margin 756000
Fixed cost 830400
Net income (loss) -74400

2) Break even unit = 830400/30 = 27680 Units

Break even sales = 27680*91 = $2518880

3) Calculate profit

Selling price Sales Profit
91 25200 25200*30-830400 = -74400
89 30200 30200*28-830400 = 15200
87 35200 35200*26-830400 = 84800
85 40200 40200*24-830400 = 134400
83 45200 45200*22-830400 = 164000
81 50200 50200*20-830400 = 173600
79 55200 55200*18-830400 = 163200

So on $81 Sales units = 50200

4) Break even unit = 830400/20 = 41520

Break even sales = 41520*81 = $3363120